OpenStudy (anonymous):

Solve: 5(x − 1) − (x + 5) = 2

OpenStudy (anonymous):

distribute that 5(x-1)

OpenStudy (anonymous):

Use foil.

OpenStudy (anonymous):

@hadleydarcym no foil here

OpenStudy (anonymous):

5x-5?

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

okay now distribute -(x+5)

OpenStudy (anonymous):

well. Okay

OpenStudy (anonymous):

-x-5?

OpenStudy (anonymous):

@hadleydarcym really @shahbaaz786 okay now add those two things together 5x-5 - x - 5

OpenStudy (anonymous):

4x-10?

OpenStudy (anonymous):

5(x - 1) - (x + 5) = 2 5x - 5 -x - 5 = 2 now combine like terms

OpenStudy (anonymous):

hoooeeekay

OpenStudy (anonymous):

yes, now set that equal to 2 4x-10 = 2 now solve for x

OpenStudy (anonymous):

good explaining robz

OpenStudy (anonymous):

x=3?

OpenStudy (anonymous):

you got it

OpenStudy (anonymous):

@kelliegirl33 thank you @shahbaaz786 correct

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you're welcome